OpenStudy (anonymous):

When three times the opposite of a number is subtracted from the number itself, the difference is six.

OpenStudy (kewlgeek555):

The opposite of what number?

OpenStudy (anonymous):

that suppose to be a variable

OpenStudy (anonymous):

3 * (-x) - x = 6 -4x = 6 x = -1.5?